# On the Analysis of Misspecified Variational Inequalities with Nonlinear Constraints

### v1

@@ 1.1 Literature Review
Saddle Point (SP) and Variational Inequality (VI) Problems. Recent research has extended the study of misspecification beyond classical optimization to SP and VI problems. In [3], a class of misspecified monotone VI problems is studied, where misspecifications may arise in the objective as well as in the constraint set. To tackle these issues, the authors develop extragradient and regularized first-order algorithms tailored for misspecified monotone VI problems coupled with strongly convex learning subproblems, thereby guaranteeing stable convergence. Complementary to this, [10] analyze stochastic variational inequality (SVI) problems in the context of misspecification, covering both stochastic convex optimization and a range of stochastic equilibrium models. Their analysis critically relies on the strong convexity of the learning problem, which guarantees convergence of both the optimization and learning iterates. They further establish almost-sure convergence in strongly monotone settings, achieving the optimal rate of $\mathcal{O}(1/K)$, and in merely monotone settings by employing iterative Tikhonov regularization, which yields a rate of $\mathcal{O}(\sqrt{\log K}/\sqrt{K})$ under weak sharpness. [3] extend their earlier work [2] to misspecified monotone variational inequality problems and show that a constant–steplength misspecified extragradient scheme achieves asymptotic convergence of the iterates. More recently, Ahmadi and Hamedani [4] examine a class of misspecified convex-concave SP problems where the misspecified parameter can be obtained through a secondary SP problem with a strongly convex-concave objective function. They introduce a Learning-aware Accelerated Primal-Dual method that integrates the parameter updates directly into the dual momentum step and utilizes an adaptive backtracking scheme to select step sizes adaptively. Their approach attains a convergence rate of $\mathcal{O}(\log K/K)$ without relying on compactness of the feasible sets. In addition, they extend their method to settings where the secondary SP problem may not have a unique solution, proving that a modified version of the algorithm tailored to this structure achieves a rate of $\mathcal{O}(1/\sqrt{K})$ in terms of both gap function and learning suboptimality.
@@ 1.2 Research Gap and Our Contribution
Table 1: Comparison of related misspecified optimization/VI studies.
@@ 1.3 Motivating Example
The Misspecified constrained VIs in problem (Misspecified VI), have broad applications in markets and networked systems where feasibility depends on misspecified parameters learned from data. Some of these examples include constrained Nash-Cournot competition characterized by demand misspecification and policy cap [9], inverse and data-driven equilibrium in [19], misspecified Markowitz portfolio optimization problem in [6], and minimax schemes with expectation constraints [22]. Next we provide a Cournot example in the following motivating example.
@@
Modern commodity markets often operate under policy rules that constrain prices while firms face uncertainty about demand responsiveness. Consider a homogeneous goods market with $N$ firms, $i=1,\dots,N$. Firm $i$ chooses a quantity $x_{i}\geq 0$; let the aggregate output be $X\triangleq\sum_{i=1}^{N}x_{i}$. The market clears at a linear inverse demand $p(X;b^{*})=a^{*}-b^{*}X,$ where $a^{*}>0$ is known and $b^{*}>0$ is the unknown slope capturing how aggressively price falls with output. Firms know costs and capacities but do not know the true $b^{*}$. Beyond firm-level capacity limits $x_{i}\in[0,\mathrm{Cap}_{i}]$, many markets also face a price-function constraint induced by policy, such as an administrative price cap. We encode this as $p(X;b^{*})\leq\delta$, for some $\delta>0$. Mathematically, it imposes a lower bound on feasible output that depends on the (unknown) slope $b^{*}$. Let $c_{i}:[0,\mathrm{Cap}_{i}]\to\mathbb{R}_{+}$ be firm $i$’s convex cost. Then each firm’s objective is to maximize the profit $\pi_{i}({\bf x})\triangleq p(X;b^{*})x_{i}-c_{i}(x_{i})$ such that $x_{i}\in[0,\mathrm{Cap}_{i}]$ and $p(X;b^{*})\leq\delta$. In practice, the firms observe sample pairs $\{(p_{\text{obs},t},X_{t})\}_{t=1}^{T}$ and must learn $b^{*}$ from data while computing equilibria that respect the price constraint.
@@
Moreover, the feasible set in (Misspecified VI) can be characterized as $\mathcal{X}(b)\triangleq\Bigl\{\mathbf{x}=[x_{i}]_{i=1}^{N}\in\prod_{i=1}^{N}[0,\text{Cap}_{i}]\Big|p(X;b^{*})\leq a^{*}-\delta\Bigr\}$. Then the operator $F(\mathbf{x},b)$ is defined by concatenating the gradients of the negative profit functions of the firms $F_{i}(\mathbf{x},b)\triangleq c_{i}^{\prime}(x_{i})+b(X+x_{i})-a^{*}$.
@@ 2 Preliminaries and Assumptions
    $\|H(\theta)-H(\vartheta)\|\leq L_{H}\|\theta-\vartheta\|^{2},\quad\left(H(\theta)-H(\vartheta)\right)^{\top}(\theta-\vartheta)\geq\mu_{H}\|\theta-\vartheta\|^{2},\quad\forall\theta,\vartheta\in\Theta.$
@@
    For any $j\in\{1,\ldots,J\}$ and $\theta\in\Theta$, $f_{j}(\cdot,\theta)$ is convex and Lipschitz continuous, with Lipschitz constants $L_{f,x}\geq 0$. Moreover, for any $x\in X$, $f_{j}(x,\cdot)$ is Lipschitz continuous, with Lipschitz constant $L_{f,\theta}\geq 0$.
@@
    The sets $X\subseteq\mathbb{R}^{n}$ and $\Theta\subseteq\mathbb{R}^{m}$ are nonempty, closed, and convex.
@@
$\nabla_{x}\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta^{*})=\sum_{j=1}^{J}\left[\rho_{k}f_{j}(x_{k},\theta^{*})+\lambda_{k}^{(j)}\right]_{+}\nabla f_{j}(x_{k},\theta^{*})$ (4)
@@
Note that any solution pair $(x^{*},\theta^{*})$ of (Misspecified VI) is a solution of the following optimization problem
@@
Consider (Misspecified VI) problem under Assumption 2.1 and let $(x^{*},\theta^{*},\lambda^{*})$ satisfy the KKT conditions of Proposition 2.1. Then for all $x\in\mathcal{X(\theta^{*})}$, $(x-x^{*})^{\top}F(x^{*},\theta^{*})+\sum_{j\in J}\lambda^{*(j)}f_{j}(x,\theta^{*})\geq 0$.
@@
The result follows directly from Lemma 1 in [5], which establishes the corresponding inequality under Assumption 2.1. ∎
@@ 3 Proposed Method
In this section, we present an augmented Lagrangian method for solving (Misspecified VI). Due to the presence of a nonlinear and misspecified constraint, we introduce a multiplier $\lambda$ to relax this constraint, following Definition 2.1. Moreover, to handle the parameter misspecification, we simultaneously solve a secondary VI problem that generates a sequence of iterates converging to the unique optimal parameter $\theta^{*}$. The approximation of $\theta^{*}$ obtained at each iteration is then used to update both the main decision variable $x$ and the multiplier $\lambda$. To evaluate the quality of the obtained solution, we introduce a relaxed gap function. First, recall that the standard VI gap function for (Misspecified VI) is defined as $\text{Gap}(x,\theta^{*})\triangleq\sup_{y\in\mathcal{X}(\theta^{*})}{F(y,\theta^{*})^{\top}(x-y)},$ and an $\epsilon$-approximate solution satisfies $\text{Gap}(x,\theta^{*})\leq\epsilon$. This gap function is a valid optimality metric, since $\text{Gap}(x,\theta^{*})\geq 0$ for any feasible $x\in\mathcal{X}(\theta^{*})$, and $\text{Gap}(x,\theta^{*})=0$ implies that $x$ is a solution to (Misspecified VI). However, because our proposed method employs a dual multiplier to enforce constraints and accounts for the misspecified parameter $\theta$, the generated sequence may be infeasible. Consequently, the algorithm’s output $\bar{x}$ may not belong to the feasible set $\mathcal{X}(\theta^{*})$, and we might have $\text{Gap}(\bar{x},\theta^{*})<0$. To address this issue, we measure both (i) the level of infeasibility, quantified as $\mathbf{1}^{\top}[f(x,\theta^{*})]\leq\epsilon$, and (ii) a relaxed gap function defined as $\widetilde{\text{Gap}}(x,\theta^{*})\triangleq\sup_{y\in\mathcal{X}_{\epsilon}(\theta^{*})}{F(y,\theta^{*})^{\top}(x-y)}$ where $\mathcal{X}_{\epsilon}(\theta^{*})\triangleq\{x\in X\mid\mathbf{1}^{\top}[f(x,\theta^{*})]\leq\epsilon\}$ denotes the $\epsilon$-enlargement of the feasible set. It is straightforward to verify that $\widetilde{\text{Gap}}(x,\theta^{*})\geq 0$ for any $x\in\mathcal{X}_{\epsilon}(\theta^{*})$, and $\widetilde{\text{Gap}}(x,\theta^{*})\geq\text{Gap}(x,\theta^{*})$ for any $x$. Therefore, we define an $\epsilon$-approximate solution $\bar{x}\in X$ if
$\text{(Relaxed gap): }\widetilde{\text{Gap}}(\bar{x},\theta^{*})\leq\epsilon,\qquad\text{(Infasibility): }\mathbf{1}^{\top}[f(\bar{x},\theta^{*})]\leq\epsilon.$
@@
To find such an approximate solution, we propose a novel single-loop algorithm by combining the forward-reflected-backward method with an augmented Lagrangian update. This combination enables us to effectively handle nonlinear constraints under misspecified evaluations of both the operator $F$ and the nonlinear constraint functions $f_{j}$’s. Recall the augmented Lagrangian function from Definition 2.1, then for fixed $(\theta,\lambda)$, the term $\Phi_{\rho}(x,\lambda,\theta)$ imposes penalization for violating the constraints $f_{j}(x,\theta)\leq 0$. In particular, at each iteration $k\geq 0$, the decision variable $x$ is updated using a projected forward operation applied to the augmented Lagrangian function combined with a reflected (correction) step $r_{k}=F(x_{k},\theta_{k})-F(x_{k-1},\theta_{k-1})$ as $x_{k+1}=\Pi_{X}[x_{k}-\gamma_{k}(r_{k}+\nabla_{x}\mathcal{L}_{\rho_{k}}(x_{k},\lambda_{k},\theta_{k}))]$. Then the dual multiplier $\lambda_{k}$ is updated via a projected gradient ascent step to promote feasibility based on the newly updated primal variable $x_{k+1}$. Finally, the misspecified parameter $\theta_{k+1}$ is refined by performing one projected gradient step associated with the secondary VI problem (2). The steps of the algorithm are outlined in Algorithm 1.
@@
1:Input: $x_{0}\in X$, $\theta_{0}\in\Theta$, $\lambda_{0}=\mathbf{0}$, $\{\rho_{k},\gamma_{k},\eta_{k}\}_{k\geq 0}\subset\mathbb{R}_{++}$.
2:for $k=0,1,2,\dots,K-1$ do
3:  $r_{k}\leftarrow F(x_{k},\theta_{k})-F(x_{k-1},\theta_{k-1})$
4:  $x_{k+1}\leftarrow\Pi_{X}\left[x_{k}-\gamma_{k}\left(F(x_{k},\theta_{k})+r_{k}+\mathbf{J}f(x_{k},\theta_{k})[\rho_{k}f(x_{k},\theta_{k})+\lambda_{k}]_{+}\right)\right]$
5:  $\lambda_{k+1}\leftarrow\left[\lambda_{k}+\rho_{k}f(x_{k+1},\theta_{k})\right]_{+}$
6:  $\theta_{k+1}\leftarrow\Pi_{\Theta}\left[\theta_{k}-\eta_{k}H(\theta_{k})\right]$
7:end for
@@ 4 Convergence Analysis
In this section, we establish the convergence properties of Algorithm 1 under Assumption 2.1. In particular, through some step-by-step analysis, we demonstrate that our method achieves a convergence rate of $\mathcal{O}(1/K)$ in terms of the relaxed gap and infeasibility metrics. All related proofs are provided in the Appendix.
We begin by showing some preliminary bounds on the difference between the constraint enforcement term of the augmented Lagrangian function, $\phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})$, and the corresponding term in the standard Lagrangian, $\sum_{j\in J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})$.
@@
$\sum_{j\in J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)\leq\frac{1}{2\rho_{k}}\left(\|\lambda_{k}-\lambda\|^{2}-\|\lambda_{k+1}-\lambda\|^{2}\right)$
$\qquad+L^{\Phi}_{\lambda\theta}\|\lambda_{k+1}-\lambda\|\|\theta_{k}-\theta^{*}\|+\rho_{k}L_{f,\theta}\|\theta_{k}-\theta^{*}\|\sum_{j\in J_{k}^{+}}|f_{j}(x_{k+1},\theta^{*})|.$
@@
$\left(x_{k+1}-x\right)^{\top}F\left(x,\theta^{*}\right)+\sum_{j\in J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\Phi_{\rho_{k}}(x,\lambda_{k},\theta^{*})$
$\quad\leq\frac{1}{2\gamma_{k}}\Big(\|x_{k}-x\|^{2}-\|x_{k+1}-x\|^{2}\Big)+\frac{1}{2\rho_{k}}\Big(\|\lambda_{k}-\lambda\|^{2}-\|\lambda_{k+1}-\lambda\|^{2}\Big)$
$\qquad+\left(\frac{L^{\Phi}+2L_{F,x}+L_{F,\theta}}{2}-\frac{1}{2\gamma_{k}}\right)\|x_{k+1}-x_{k}\|^{2}+\langle r_{k+1},x_{k+1}-x\rangle-\langle r_{k},x_{k}-x\rangle$
$\qquad+\frac{L_{F,\theta}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}+L_{F,\theta}\|\theta_{k+1}-\theta^{*}\|\|x_{k+1}-x\|+L^{\Phi}_{x\theta}\|\theta_{k}-\theta^{*}\|\Big(\|x_{k+1}-x_{k}\|+\|x_{k}-x\|\Big)$
$\qquad+L_{\lambda\theta}^{\Phi}\|\lambda_{k+1}-\lambda\|\|\theta_{k}-\theta^{*}\|+\rho_{k}L_{f,\theta}\|\theta_{k}-\theta^{*}\|\sum_{j\in J_{k}^{+}}|f_{j}(x_{k+1},\theta^{*})|,$
@@
See Appendix B for the proof. ∎
Note that the second-to-last term on the right-hand side of the inequality contains the product $\|\lambda_{k+1}-\lambda\|\theta_{k}-\theta^{*}\|$. Assuming that the learning iterates $\{\theta_{k}\}$ are converging to $\theta^{*}$, this term will be summable only if the dual iterates do not grow. In fact, in the following lemma, we establish that the dual iterates $\{\lambda_{k}\}_{k\geq 0}$ are bounded.

### v3

@@ 1.1 Literature Review
Saddle Point (SP) and Variational Inequality (VI) Problems. Recent research has extended the study of misspecification beyond classical optimization to SP and VI problems. In [3], a class of misspecified monotone VI problems is studied, where misspecifications may arise in the objective. To tackle these issues, the authors develop extragradient and regularized first-order algorithms tailored for misspecified monotone VI problems coupled with strongly convex learning subproblems, thereby guaranteeing stable convergence. Complementary to this, [10] analyze stochastic variational inequality (SVI) problems in the context of misspecification, covering both stochastic convex optimization and a range of stochastic equilibrium models. Their analysis critically relies on the strong convexity of the learning problem, which guarantees convergence of both the optimization and learning iterates. They further establish almost-sure convergence in strongly monotone settings, achieving the optimal rate of $\mathcal{O}(1/K)$, and in merely monotone settings by employing iterative Tikhonov regularization, which yields a rate of $\mathcal{O}(\sqrt{\log K}/\sqrt{K})$ under weak sharpness. [3] extend their earlier work [2] to misspecified monotone variational inequality problems and show that a constant–steplength misspecified extragradient scheme achieves asymptotic convergence of the iterates. More recently, Ahmadi and Hamedani [4] examine a class of misspecified convex-concave SP problems where the misspecified parameter can be obtained through a secondary SP problem with a strongly convex-concave objective function. They introduce a Learning-aware Accelerated Primal-Dual method that integrates the parameter updates directly into the dual momentum step and utilizes an adaptive backtracking scheme to select step sizes adaptively. Their approach attains a convergence rate of $\mathcal{O}(\log K/K)$ without relying on compactness of the feasible sets. In addition, they extend their method to settings where the secondary SP problem may not have a unique solution, proving that a modified version of the algorithm tailored to this structure achieves a rate of $\mathcal{O}(1/\sqrt{K})$ in terms of both gap function and learning suboptimality.
@@ 1.2 Research Gap and Our Contribution
Table 1: Comparison of related misspecified convex optimization/monotone VI studies.
@@ 1.3 Motivating Example
The Misspecified constrained VIs in problem (Misspecified VI) have broad applications in markets and networked systems where feasibility depends on misspecified parameters learned from data. Some of these examples include constrained Nash-Cournot competition characterized by demand misspecification and policy cap [9], inverse and data-driven equilibrium in [19], misspecified Markowitz portfolio optimization problem in [6], and minimax schemes with expectation constraints [22]. Next, we provide a Cournot example in the following motivating example.
@@
Modern commodity markets often operate under policy rules that constrain prices while firms face uncertainty about demand responsiveness. Consider a homogeneous goods market with $N$ firms, $i=1,\dots,N$. Firm $i$ chooses a quantity $x_{i}\geq 0$; let the aggregate output be $X\triangleq\sum_{i=1}^{N}x_{i}$. The market clears at a linear inverse demand $p(X;b^{*})=a^{*}-b^{*}X,$ where $a^{*}>0$ is known and $b^{*}>0$ is the unknown slope capturing how aggressively price falls with output. Firms know costs and capacities but do not know the true $b^{*}$. Beyond firm-level capacity limits $x_{i}\in[0,\mathrm{Cap}_{i}]$, many markets also face a price-function constraint induced by policy, such as an administrative price cap. We encode this as $p(X;b^{*})\leq a^{*}-\delta$, for some $\delta>0$. Mathematically, it imposes a lower bound on the feasible output that depends on the (unknown) slope $b^{*}$. Let $c_{i}:[0,\mathrm{Cap}_{i}]\to\mathbb{R}_{+}$ be firm $i$’s convex cost. Then, each firm’s objective is to maximize the profit $\pi_{i}({\bf x})\triangleq p(X;b^{*})x_{i}-c_{i}(x_{i})$ such that $x_{i}\in[0,\mathrm{Cap}_{i}]$ and $p(X;b^{*})\leq a^{*}-\delta$. In practice, the firms observe sample pairs $\{(p_{\text{obs},t},X_{t})\}_{t=1}^{T}$ and must learn $b^{*}$ from data while computing equilibria that respect the price constraint.
@@
Moreover, the feasible set in (Misspecified VI) can be characterized as $\mathcal{X}(b^{*})\triangleq\Bigl\{\mathbf{x}=[x_{i}]_{i=1}^{N}\in\prod_{i=1}^{N}[0,\text{Cap}_{i}]\Big|p(X;b^{*})\leq a^{*}-\delta\Bigr\}$. Then, the operator $F(\mathbf{x},b)$ is defined by concatenating the gradients of the negative profit functions of the firms $F_{i}(\mathbf{x},b)\triangleq c_{i}^{\prime}(x_{i})+b(X+x_{i})-a^{*}$.
@@ 2 Preliminaries and Assumptions
    $\|H(\theta)-H(\vartheta)\|\leq L_{H}\|\theta-\vartheta\|,\quad\left(H(\theta)-H(\vartheta)\right)^{\top}(\theta-\vartheta)\geq\mu_{H}\|\theta-\vartheta\|,\quad\forall\theta,\vartheta\in\Theta.$
@@
    For any $j\in\{1,\ldots,J\}$ and $\theta\in\Theta$, the function $f_{j}(\cdot,\theta)$ is convex and continuously differentiable on $X$. Moreover, there exists $L_{\nabla f,x}\geq 0$ such that for all $x,y\in X$ and all $\theta\in\Theta$,
    $\|\nabla_{x}f_{j}(x,\theta)-\nabla_{x}f_{j}(y,\theta)\|\leq L_{\nabla f,x}\|x-y\|.$
    In addition, for any $\theta\in\Theta$ and $x\in X$, $f_{j}(\cdot,\theta)$ and $f_{j}(x,\cdot)$ are Lipschitz continuous with constants $L_{f,x}\geq 0$ and $L_{f,\theta}\geq 0$ respectively.
@@
    The sets $X\subseteq\mathbb{R}^{n}$ and $\Theta\subseteq\mathbb{R}^{m}$ are nonempty, closed, convex and compact.
@@
$\nabla_{x}\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta^{*})=\sum_{j=1}^{J}\left[\rho_{k}f_{j}(x_{k},\theta^{*})+\lambda_{k}^{(j)}\right]_{+}\nabla f_{j}(x_{k},\theta^{*}).$ (4)
@@
Proof. Note that any solution pair $(x^{*},\theta^{*})$ of (Misspecified VI) is a solution of the following optimization problem
@@
Consider (Misspecified VI) problem under Assumption 2.1 and let $(x^{*},\theta^{*},\lambda^{*})$ satisfy the KKT conditions of Proposition 2.1. Then for all $x\in\mathcal{X(\theta^{*})}$, $(x-x^{*})^{\top}F(x^{*},\theta^{*})+\sum_{j=1}^{J}\lambda^{*(j)}f_{j}(x,\theta^{*})\geq 0$.
@@
Proof The result follows directly from Lemma 1 in [5], which establishes the corresponding inequality under Assumption 2.1. ∎
@@ 3 Proposed Method
In this section, we present an augmented Lagrangian method for solving (Misspecified VI). Due to the presence of a nonlinear misspecified constraint, we introduce a multiplier $\lambda$ to relax this constraint, following Definition 2.1. Moreover, to handle the parameter misspecification, we simultaneously solve a secondary VI problem that generates a sequence of iterates converging to the unique optimal parameter $\theta^{*}$. The approximation of $\theta^{*}$ obtained at each iteration is then used to update both the main decision variable $x$ and the multiplier $\lambda$. To evaluate the quality of the obtained solution, we introduce a relaxed gap function. First, recall that the standard VI gap function for (Misspecified VI) is defined as $\text{Gap}(x,\theta^{*})\triangleq\sup_{y\in\mathcal{X}(\theta^{*})}{F(y,\theta^{*})^{\top}(x-y)},$ and an $\epsilon$-approximate solution satisfies $\text{Gap}(x,\theta^{*})\leq\epsilon$. This gap function is a valid optimality metric, since $\text{Gap}(x,\theta^{*})\geq 0$ for any feasible $x\in\mathcal{X}(\theta^{*})$, and $\text{Gap}(x,\theta^{*})=0$ implies that $x$ is a solution to (Misspecified VI). However, because our proposed method employs a dual multiplier to enforce constraints and accounts for the misspecified parameter $\theta$, the generated sequence may be infeasible. Consequently, the algorithm’s output $\bar{x}$ may not belong to the feasible set $\mathcal{X}(\theta^{*})$, and we might have $\text{Gap}(\bar{x},\theta^{*})<0$. To address this issue, we measure both (i) the level of infeasibility, quantified as $\mathbf{1}^{\top}[f(x,\theta^{*})]_{+}\leq\epsilon$, and (ii) a relaxed gap function defined as $\widetilde{\text{Gap}}(x,\theta^{*})\triangleq\sup_{y\in\mathcal{X}_{\epsilon}(\theta^{*})}{F(y,\theta^{*})^{\top}(x-y)}$ where $\mathcal{X}_{\epsilon}(\theta^{*})\triangleq\{x\in X\mid\mathbf{1}^{\top}[f(x,\theta^{*})]_{+}\leq\epsilon\}$ denotes the $\epsilon$-enlargement of the feasible set. It is straightforward to verify that $\widetilde{\text{Gap}}(x,\theta^{*})\geq 0$ for any $x\in\mathcal{X}_{\epsilon}(\theta^{*})$, and $\widetilde{\text{Gap}}(x,\theta^{*})\geq\text{Gap}(x,\theta^{*})$ for any $x$. Therefore, we define an $\epsilon$-approximate solution $\bar{x}\in X$ if
$\text{(Relaxed gap): }\widetilde{\text{Gap}}(\bar{x},\theta^{*})\leq\epsilon,\qquad\text{(Infasibility): }\mathbf{1}^{\top}[f(\bar{x},\theta^{*})]_{+}\leq\epsilon.$
@@
To find such an approximate solution, we propose a novel single-loop algorithm by combining the forward-reflected-backward method with an augmented Lagrangian update. This combination enables us to effectively handle nonlinear constraints under misspecified evaluations of both the operator $F$ and the nonlinear constraint functions $f_{j}$’s. Recall the augmented Lagrangian function from Definition 2.1, then for fixed $(\theta,\lambda)$, the term $\Phi_{\rho}(x,\lambda,\theta)$ imposes penalization for violating the constraints $f_{j}(x,\theta)\leq 0$. In particular, at each iteration $k\geq 0$, the decision variable $x$ is updated using a projected forward operation applied to the augmented Lagrangian function combined with a reflected (correction) step $r_{k}=F(x_{k},\theta_{k})-F(x_{k-1},\theta_{k-1})$ as $x_{k+1}=\Pi_{X}[x_{k}-\gamma_{k}(r_{k}+\nabla_{x}\mathcal{L}_{\rho_{k}}(x_{k},\lambda_{k},\theta_{k}))]$. Then, the dual multiplier $\lambda_{k}$ is updated via a projected gradient ascent step to promote feasibility based on the newly updated primal variable $x_{k+1}$. Finally, the misspecified parameter $\theta_{k+1}$ is refined by performing one projected gradient step associated with the secondary VI problem (2). The steps of the algorithm are outlined in Algorithm 1.
@@
1:Input: $x_{0}\in X$, $\theta_{0}\in\Theta$, $\{\rho_{k},\gamma_{k},\eta_{k}\}_{k\geq 0}\subset\mathbb{R}_{++}$.
2:Initialization: $x_{-1}\leftarrow x_{0}$, $\theta_{-1}\leftarrow\theta_{0}$, $\lambda_{0}\leftarrow\mathbf{0}$
3:for $k=0,1,2,\dots,K-1$ do
4:  $r_{k}\leftarrow F(x_{k},\theta_{k})-F(x_{k-1},\theta_{k-1})$
5:  $x_{k+1}\leftarrow\Pi_{X}\left[x_{k}-\gamma_{k}\left(F(x_{k},\theta_{k})+r_{k}+\mathbf{J}f(x_{k},\theta_{k})[\rho_{k}f(x_{k},\theta_{k})+\lambda_{k}]_{+}\right)\right]$
6:  $\lambda_{k+1}\leftarrow\left[\lambda_{k}+\rho_{k}f(x_{k+1},\theta_{k})\right]_{+}$
7:  $\theta_{k+1}\leftarrow\Pi_{\Theta}\left[\theta_{k}-\eta_{k}H(\theta_{k})\right]$
8:end for
@@ 4 Convergence Analysis
In this section, we establish the convergence properties of Algorithm 1 under Assumption 2.1. In particular, through a step-by-step analysis, we demonstrate that our method achieves a convergence rate of $\mathcal{O}(1/K)$ in terms of the relaxed gap and infeasibility metrics. All related proofs are provided in the Appendix.
We begin by showing some preliminary bounds on the difference between the constraint enforcement term of the augmented Lagrangian function, $\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})$, and the corresponding term in the standard Lagrangian, $\sum^{J}_{j=1}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})$.
@@
$-\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)+\sum_{j=1}^{J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})\leq\frac{1}{2\rho_{k}}\left(\|\lambda_{k}-\lambda\|^{2}-\|\lambda_{k+1}-\lambda\|^{2}\right)+\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}$
$\qquad+L^{\Phi}_{\lambda\theta}\|\lambda_{k+1}-\lambda\|\|\theta_{k}-\theta^{*}\|+2L_{f,\theta}(J\rho_{k}D_{f}+\sqrt{J}\|\lambda_{k}\|)\|\theta_{k}-\theta^{*}\|.$
@@
Lemma 4.2 (Lipschitz continuity of $\nabla_{x}\Phi_{\rho}(\cdot,\lambda,\theta)$).
Suppose Assumption 2.1 holds. Then, for every $\rho>0$, every $\lambda\in\mathbb{R}^{J}$, and every $\theta\in\Theta$, the mapping $\nabla_{x}\Phi_{\rho}(\cdot,\lambda,\theta)$ is Lipschitz continuous on $X$. In particular, for any $x,y\in X$, $\|\nabla_{x}\Phi_{\rho}(x,\lambda,\theta)-\nabla_{x}\Phi_{\rho}(y,\lambda,\theta)\|\leq(\rho C_{1}+\|\lambda\|C_{2})\|x-y\|,$ where $C_{1}\triangleq\sqrt{J}\big(L_{\nabla f,x}D_{f}+L_{f,x}M_{\nabla f}\big),\quad C_{2}\triangleq\sqrt{J}\,L_{\nabla f,x},$ and $D_{f}\triangleq\sup_{(z,\vartheta)\in X\times\Theta}\|[f(z,\vartheta)]_{+}\|<\infty,\quad M_{\nabla f}\triangleq\sup_{(z,\vartheta)\in X\times\Theta}\|\nabla_{x}f(z,\vartheta)\|<\infty.$
Proof.
Proof See Appendix B for the proof. ∎
@@
$\left(x_{k+1}-x\right)^{\top}F\left(x,\theta^{*}\right)+\sum_{j=1}^{J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\Phi_{\rho_{k}}(x,\lambda_{k},\theta^{*})$
$\quad\leq\frac{1}{2\gamma_{k}}\Big(\|x_{k}-x\|^{2}-\|x_{k+1}-x\|^{2}\Big)+\frac{1}{2\rho_{k}}\Big(\|\lambda_{k}-\lambda\|^{2}-\|\lambda_{k+1}-\lambda\|^{2}\Big)+\frac{L_{F,\theta}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}$
$\qquad+\left(\frac{\rho_{k}C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}}{2}-\frac{1}{2\gamma_{k}}\right)\|x_{k+1}-x_{k}\|^{2}+\frac{C_{2}}{2}\|\lambda_{k}-\lambda^{*}\|\|x_{k+1}-x_{k}\|^{2}$
$\qquad+\langle r_{k+1},x_{k+1}-x\rangle-\langle r_{k},x_{k}-x\rangle+L_{F,\theta}\|\theta_{k+1}-\theta^{*}\|\|x_{k+1}-x\|+\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}$
$\qquad+L^{\Phi}_{x\theta}\|\theta_{k}-\theta^{*}\|\Big(\|x_{k+1}-x_{k}\|+\|x_{k}-x\|\Big)+\frac{L_{F,x}}{2}\left(\|x_{k}-x_{k-1}\|^{2}-\|x_{k+1}-x_{k}\|^{2}\right)$
$\qquad+\frac{L_{F,\theta}}{2}\left(\|\theta_{k}-\theta_{k-1}\|^{2}-\|\theta_{k+1}-\theta_{k}\|^{2}\right)+L_{\lambda\theta}^{\Phi}\|\lambda_{k+1}-\lambda\|\|\theta_{k}-\theta^{*}\|$
$\qquad+2L_{f,\theta}(J\rho_{k}D_{f}+\sqrt{J}\|\lambda_{k}\|)\|\theta_{k}-\theta^{*}\|,$
@@
Proof See Appendix C for the proof. ∎
Note that the last term on the right-hand side of the inequality contains the product $\|\lambda_{k+1}-\lambda\|\|\theta_{k}-\theta^{*}\|$. Assuming that the learning iterates $\{\theta_{k}\}$ are converging to $\theta^{*}$, this term will be summable only if the dual iterates do not grow. In fact, in the following lemma, we establish that the dual iterates $\{\lambda_{k}\}_{k\geq 0}$ are bounded.